# Learning for Dynamic Combinatorial Optimization without Training Data

### v1

@@ 4.2 DyCO-GNN: fast convergence and robust solution quality
To address the shortcomings of naively warm-starting PI-GNN, we propose DyCO-GNN—a simple yet effective method that facilitates both fast convergence and robust solution quality across graph snapshots. Unlike standard warm start, which reuses the exact model parameters from the previous snapshot, DyCO-GNN integrates a strategic initialization method—shrink and perturb (SP) [2]. The goal is to retain the benefits of accelerated convergence while mitigating the tendency of the model to become trapped in suboptimal local minima due to overconfident predictions caused by warm start.
@@
When a new snapshot $G_{p}^{(t)}$ arrives, DyCO-GNN applies SP to the warm-started parameters before starting the optimization process. The newly initialized parameters are defined as
$\theta^{(t)}\leftarrow\lambda_{\text{shrink}}\theta^{(t-1)}+\lambda_{\text{perturb}}\epsilon^{(t)},$
where $10<\lambda_{\text{shrink}}<$ 1, $10<\lambda_{\text{perturb}}<$ 1, and $\epsilon^{(t)}\sim\mathcal{N}(0,\sigma^{2})$. With the reintroduced gradient diversity and disrupted premature overconfidence, DyCO-GNN effectively escapes the local minima in the loss landscape. This gentle destabilization promotes exploration of alternative descent paths, functioning as a principled soft reset mechanism that facilitates more effective and robust optimization.
With this simple design, DyCO-GNN achieves a balanced trade-off: it preserves the efficiency of warm start under tight time constraints while enabling the model to reach higher-quality solutions as more computational budget becomes available. The pseudocode of DyCO-GNN is provided in Algorithm 2. As we show in Section 5, DyCO-GNN consistently outperforms warm-started PI-GNN across DCO problems such as dynamic MaxCut, MIS, and TSP, particularly under moderate to generous time budgets where naive warm start fails to find high-quality solutions effectively. We will also show that DyCO-GNN finds better solutions than the converged static PI-GNN much faster.
@@ 4.3 Theoretical support
We analytically support the advantage of SP over warm start. The theorem is based on the Goemans and Williamson [13] (GW) algorithm, the best-known optimization algorithm for solving MaxCut. The GW algorithm is structured similar to DyCO-GNN, but with a more computationally intensive QUBO relaxation (an SDP) and rounding step (randomized projections). We adopt it to show that the advantage of SP persists even if we advance the relaxation and rounding steps.
@@
In words, the theorem shows that (pre-rounding) perturbations can strictly increase the probability of finding the optimal cut. The proof is shown in Appendix D, along with a detailed description of the GW algorithm, and a corollary extending the result to perturbations of the GW SDP initialization.
@@ 5 Experiments
We empirically evaluate DyCO-GNN on dynamic instances of MaxCut, MIS, and TSP of varying problem sizes. Baselines include static PI-GNN and warm-started PI-GNN. We exclude methods that rely on extensive offline training across instance distributions. This choice aligns with the core design philosophy of our proposed approach, which, like PI-GNN, emphasizes instance-specific adaptability over generalization. While generalizable methods trained on large datasets may offer strong performance on similar distributions, prior research has shown that methods from the PI-GNN family often achieve comparable performance across static CO problems like MaxCut and MIS [42, 51, 19]. Moreover, the generalization gap of the methods with generalizability increases significantly when distribution shift is present [25, 51]. By focusing our comparison on methods within the PI-GNN family, we provide a clearer assessment of the algorithmic innovations without conflating results with the (dis)advantages of dataset-level generalization.
@@ 5.1 Datasets
We utilized established graph datasets that contain temporal information to ensure that the dynamic evolution of the graphs accurately reflects realistic, time-dependent structural changes. This approach allows us to capture authentic patterns of change within the graph over time, thereby enhancing the validity and applicability of our experimental results. We summarize the statistics of the original and preprocessed datasets in Table 1.
Original Preprocessed
Dataset Nodes Edges Nodes Edges $\Delta$ edges
Infectious 410 17,298 410 2765 $\sim$276 (10%)
UC Social 1899 59,835 1899 13,838 $\sim$1384 (10%)
DBLP 12,590 49,759 12,590 49,636 $\sim$993 (2%)
burma14 14 91 15 105 -
ulysses22 22 213 23 253 -
st70 70 2,415 71 2,485 -
For MaxCut and MIS, we employed Infectious [21], a human contact network; UC Social [36], a communication network; and DBLP [31], a citation network. All datasets are publicly available in the Koblenz Network Collection [30]. Each dataset is a graph with timestamps attached to all edges. For preprocessing, we began by sorting all edge events in chronological order, followed by the removal of self-loops and duplicate edges. Finally, we made all graphs undirected. To construct DTDG snapshots for MaxCut, we converted the graphs into 10 snapshots by linearly increasing the number of edges to include in each of them (i.e., the dynamic graph grows over time). For Infectious and UC Social, we added $\sim$10of total edges of the final graph every step; for DBLP, we added $\sim$2each step, considering its total number of edges is much larger. Edge additions will lead to MIS constraint violations. In that case, a postprocessing step that reuses previous solutions and removes violations would be sufficient to solve the DCO problem. Therefore, to construct DTDG snapshots for MIS, we reversed the snapshot order and turn edge additions into edge deletions.
For TSP, we first took static TSP benchmark problems from TSPLIB [40]. Specifically, we considered burma14, ulysses22, and st70. We added an extra node and let it move along a straight line within the region defined by the other existing nodes. We recorded the coordinates of 5 equally spaced locations along the trajectory (i.e., 5 snapshots in total).
@@ 5.2 Implementation details
DyCO-GNN consists of a node embedding layer and two graph convolution layers. Unlike previous works [42, 17, 19] that used different embedding and intermediate hidden dimensions for each problem instance, we used 512 for the embedding dimension and 256 for the hidden dimension in all our experiments. We experimented with the graph convolution operator (GCNConv) from [28] and the GraphSAGE operator (SAGEConv) from [15]. We found that DyCO-GNN with GCNConv failed to find good enough solutions to TSP, so we only report results of the SAGEConv models on dynamic TSP. We set $\lambda_{\text{shrink}}=0.4$ and $\lambda_{\text{perturb}}=0.1$ without further tuning. All MaxCut and MIS experiments were repeated five times; all TSP experiments were repeated ten times. We obtained the ground truth for each snapshot by formulating it as a QUBO problem and solving it using the Gurobi solver [14] with a time limit of 60 seconds. Additional implementation details are provided in Appendix B.
@@ 5.3 Empirical results on DCO problems
Tables 2, 3, and 4 report the mean ApR of all methods on dynamic MaxCut, MIS, and TSP, respectively. We also evaluate variants of DyCO-GNN that apply SP to different layers of DyCO-GNN. For MaxCut and MIS, results are reported using the final model checkpoint. For TSP, both the final and the best-performing checkpoints are considered. The wall-clock time of the best-performing checkpoint includes the total decoding time over all checkpoints available. As noted in Section 4, the advantage of warm-started PI-GNN compared to static PI-GNN diminishes quickly when we relax the runtime constraint. DyCO-GNN closes this performance gap and consistently outperforms static PI-GNN and warm-started PI-GNN across all problems and datasets. Even on a strict budget (first column of each table), DyCO-GNN surpasses warm-started PI-GNN in most cases. Crucially, DyCO-GNN consistently finds better solutions than fully converged static PI-GNN, and often within just 1.67% to 33.33% of the total runtime. This improvement possibly stems from information carryover and continual learning on a similar graph structure, enabled by the utilization of the previously learned hypothesis. Although the three versions of DyCO-GNN give strong results, no single setting dominates all tasks and datasets. We will discuss the implications in Section 6. We visualize snapshot-level ApRs under different time budgets in Figures 2, 3, and 4. Notably, DyCO-GNN maintains its edge across almost all evaluated snapshots, confirming its robustness. Additional experimental results are provided in Appendix C.
@@
Method 500/0.43 1000/0.85 2000/1.70 3000/2.54 5000/4.23 10000/8.46 20000/16.95
st70 PI-GNN (static) 2.01621 1.98448 1.92607 1.82142 1.59433 1.43824 1.36945
PI-GNN (warm) 1.47258 1.46055 1.42443 1.42589 1.38696 1.36187 1.34213
DyCO-GNN (emb) 1.45563 1.42486 1.38436 1.36485 1.35266 1.30753 1.29239
DyCO-GNN (GNN) 1.53036 1.44924 1.36227 1.33620 1.32063 1.27857 1.23655
DyCO-GNN (full) 1.56795 1.54532 1.38231 1.34845 1.30566 1.27127 1.24412
Sensitivity of different methods to varying degrees of change. Tables 2 and 3 show that the quality gap of converged solutions between static PI-GNN and DyCO-GNN (resp. warm-started PI-GNN) is larger (resp. smaller) for DBLP. This is because we constructed the DTDG for DBLP with a smaller relative change in the edge set ($\sim$2of total edges), which leads to greater structural overlaps. Consequently, warm-started PI-GNN and DyCO-GNN would be stronger in such a scenario. We further conducted a sensitivity analysis by varying the degrees of change in the DTDGs. Results are illustrated in Figure 5. When $\Delta\text{edges}$ is getting smaller, the gap between static and warm-started PI-GNN is narrowed, and the advantage of DyCO-GNN compared to static PI-GNN is more salient, which is consistent with our hypothesis.
@@ 6 Discussion
Limitations and future work. In our experiments, most hyperparameters were either kept at their default settings or chosen with minimal tuning. The performance of DyCO-GNN can likely be further improved by tuning the model architecture and optimizer using the snapshot(s) available at hand, and incorporating techniques such as learning rate scheduling. We have shown that applying SP to different layers of DyCO-GNN yields varying performance. A promising direction for future work is to make the SP step adaptive. Currently, the embedding layer and the GNN layers are jointly optimized without explicit decoupling. By introducing appropriate regularization strategies or auxiliary loss functions, we can promote distinct functional roles across layers—for instance, encouraging the embedding layer to capture representations specific to both the problem and graph structure and the GNN layers to learn to aggregate such representations. Additionally, if we evaluate the likelihood of changes in node assignments under dynamic conditions based on certain heuristics, it becomes possible to apply SP in a node-wise, adaptive manner. Another promising extension is to combine DyCO-GNN with the continuous relaxation annealing (CRA) technique proposed in [19] to enhance the solution quality, particularly on dense graphs.
Conclusion. We introduced DyCO-GNN, the first learning-based framework designed to solve DCO problems. Through extensive experiments on dynamic MaxCut, MIS, and TSP, we demonstrated that DyCO-GNN consistently outperforms both static and warm-started PI-GNN baselines under varying runtime constraints and across diverse problem settings. Our analysis reveals that DyCO-GNN achieves superior solution quality even under strict time budgets, always surpassing the best-performing checkpoints of the baseline methods in a fraction of their runtime. These improvements are significant in rapidly changing environments, where timely decision-making is critical. Furthermore, DyCO-GNN’s ability to efficiently adapt to evolving problem instance snapshots without optimization from scratch or external supervision underscores its potential for real-world deployment in dynamic, resource-constrained settings.
@@ Appendix A Additional related work
DCO problems have been addressed from a theoretical perspective [35, 46], including for the dynamic MaxCut [52], MIS [3], and TSP [4] problems, as well as from the perspective of designing heuristics [53]. Our work brings a learning-based approach to this line of work.
The idea of warm-starting semidefinite programs in general, and the starting solutions of interior-point methods (which can be used to solved SDPs, among other problem classes) has been explored from a theoretical perspective [54, 22, 1]. Our theoretical findings complement this literature by exploring the impacts of warm-starting SDP solutions for the MaxCut problem. Our learning-based method DyCO-GNN also brings similar ideas into the realm of learning-based DCO methods.
Appendix B Additional implementation details
We used the Adam optimizer [27]; the learning rate was set to 0.001 for all MaxCut and MIS experiments, 0.0002 for burma14 and st70, and 0.0005 for ulysses22. As described in Algorithms 1 and 2, we optimize over the first snapshot for $\text{epoch}_{\text{max}}$ epochs. Thus, we skip the first snapshot during evaluation since the results will be identical across all methods. We set $\text{epoch}_{\text{max}}=3000$ for all MaxCut and MIS experiments, $\text{epoch}_{\text{max}}=10000$ for burma14 and ulysses22, and $\text{epoch}_{\text{max}}=20000$ for st70.
The penalty term $M$ was set to 2 for MIS and $2\times\max_{(i,j)\in E}w_{ij}$ (i.e., 2×2$\times$2 ×the max distance between any two nodes) for TSP. For MIS, we postprocessed the output of our model by greedily removing violations. For TSP, we decoded the route step by step. More specifically, for burma14 and ulysses22, we discarded nodes that have already been visited and selected the most likely node to visit based on our model output; for st70, we found that this greedy decoding failed to find good enough routes and instead applied beam search that expands the top 5 possible valid nodes at each step.
All models were implemented using PyTorch [38] and PyTorch Geometric [9]. Experiments were conducted on a machine with a single NVIDIA GeForce RTX 4090 GPU, a 32-core Intel Core i9-14900K CPU, and 64 GB of RAM running Ubuntu 24.04.
Appendix C Additional experimental results
C.1 DyCO-GNN with SAGEConv for dynamic Maxcut and MIS

### v2

@@ 4.2 DyCO-GNN: fast convergence and robust solution quality
To address the shortcomings of naively warm-starting PI-GNN, we propose DyCO-GNN, a simple yet effective method that facilitates both fast convergence and robust solution quality across graph snapshots. Unlike standard warm start, which reuses the exact model parameters from the previous snapshot, DyCO-GNN integrates a strategic initialization method, shrink and perturb (SP), originally proposed in (Ash and Adams, 2020) in a different context for supervised learning tasks. The goal is to retain the benefits of accelerated convergence while mitigating the tendency of the model to become trapped in suboptimal local minima due to overconfident predictions caused by warm start.
@@
When a new snapshot $G_{p}^{t}$ arrives, DyCO-GNN applies SP to the warm-started parameters before starting the optimization process. The newly initialized parameters are defined as
$\theta^{t}\leftarrow\lambda_{\text{shrink}}\theta^{t-1}+\lambda_{\text{perturb}}\epsilon^{t},$
where $0<\lambda_{\text{shrink}}<1$, $0<\lambda_{\text{perturb}}<1$, and $\epsilon^{t}\sim\mathcal{N}(0,\sigma^{2})$. With the reintroduced gradient diversity and disrupted premature overconfidence, DyCO-GNN effectively escapes the local minima in the loss landscape. This gentle destabilization promotes exploration of alternative descent paths, functioning as a principled soft reset mechanism that facilitates more effective and robust optimization.
With this simple design, DyCO-GNN achieves a balanced trade-off: it preserves the efficiency of warm start under tight time constraints while enabling the model to reach higher-quality solutions as more computational budget becomes available. The pseudocode of DyCO-GNN is provided in Algorithm 2. As we show in Section 5, DyCO-GNN consistently outperforms warm-started PI-GNN across DCO problems, including dynamic MaxCut, MIS, and TSP, particularly under moderate to generous time budgets where naive warm start fails to find high-quality solutions effectively. We will also show that DyCO-GNN finds better solutions than the converged static PI-GNN much faster.
@@ 4.3 Theoretical support
We analytically support the advantage of SP over warm start. The theorem is based on the Goemans and Williamson (1995) (GW) algorithm, the best-known optimization algorithm for solving MaxCut. The GW algorithm is structured similar to DyCO-GNN, but with a more computationally intensive QUBO relaxation (an SDP) and rounding step (randomized projections). We adopt it to show that the advantage of SP persists even if we advance the relaxation and rounding steps.
@@
In words, the theorem shows that (pre-rounding) perturbations can strictly increase the probability of finding the optimal cut. The proof is shown in Appendix E, along with a detailed description of the GW algorithm, and a corollary extending the result to perturbations of the GW SDP initialization.
@@ 5 Experiments
We empirically evaluate DyCO-GNN on dynamic instances of MaxCut, MIS, and TSP of varying problem sizes. Baselines include static PI-GNN and warm-started PI-GNN. We exclude methods that rely on extensive offline training across instance distributions. This choice aligns with the core design philosophy of our proposed approach, which, like PI-GNN, emphasizes instance-specific adaptability over generalization. While generalizable methods trained on large datasets may offer strong performance on similar distributions, prior research has shown that methods from the PI-GNN family often achieve comparable performance across static CO problems like MaxCut and MIS (Schuetz et al., 2022; Wang and Li, 2023; Ichikawa, 2024). Moreover, the generalization gap of the methods with generalizability increases significantly when distribution shift is present (Karalias and Loukas, 2020; Wang and Li, 2023). By focusing our comparison on methods within the PI-GNN family, we provide a clearer assessment of the algorithmic innovations without conflating results with the (dis)advantages of dataset-level generalization. We also compare with the results of non-neural baselines in Appendix D.3.
@@ 5.1 Datasets
We utilized established graph datasets that contain temporal information to ensure that the dynamic evolution of the graphs accurately reflects realistic, time-dependent structural changes. This approach allows us to capture authentic patterns of change within the graph over time, thereby enhancing the validity and applicability of our experimental results. We summarize the statistics of the original and preprocessed datasets in Table 4 in Appendix B.
For MaxCut and MIS, we employed Infectious (Isella et al., 2011), a human contact network; UC Social (Opsahl and Panzarasa, 2009), a communication network; and DBLP (Ley, 2002), a citation network. All datasets are publicly available in the Koblenz Network Collection (Kunegis, 2013). Each dataset is a graph with timestamps attached to all edges. For preprocessing, we began by sorting all edge events in chronological order, followed by the removal of self-loops and duplicate edges. Finally, we made all graphs undirected. To construct DTDG snapshots for MaxCut, we converted the graphs into 10 snapshots by linearly increasing the number of edges to include in each of them (i.e., the dynamic graph grows over time). For Infectious and UC Social, we added $\sim$10% of total edges of the final graph every step; for DBLP, we added $\sim$2% each step, considering its total number of edges is much larger. Edge additions will lead to MIS constraint violations. In that case, a postprocessing step that reuses previous solutions and removes violations would be sufficient to solve the DCO problem. Therefore, to construct DTDG snapshots for MIS, we reversed the snapshot order and turn edge additions into edge deletions.
For TSP, we first took static TSP benchmark problems from TSPLIB (Reinelt, 1991). Specifically, we considered burma14, ulysses22, and st70. We added an extra node and let it move along a straight line within the region defined by the other existing nodes. We recorded the coordinates of 5 equally spaced locations along the trajectory (i.e., 5 snapshots in total).
@@ 5.2 Implementation details
DyCO-GNN consists of a node embedding layer and two graph convolution layers. Unlike previous works (Schuetz et al., 2022; Heydaribeni et al., 2024; Ichikawa, 2024) that used different embedding and intermediate hidden dimensions for each problem instance, we used 512 for the embedding dimension and 256 for the hidden dimension in all our experiments. We experimented with the graph convolution operator (GCNConv) from Kipf and Welling (2017) and the GraphSAGE operator (SAGEConv) from Hamilton et al. (2017). We found that DyCO-GNN with GCNConv failed to find good enough solutions to TSP, so we only report results of the SAGEConv models on dynamic TSP. We set $\lambda_{\text{shrink}}=0.4$ and $\lambda_{\text{perturb}}=0.1$ without further tuning. All MaxCut and MIS experiments were repeated five times; all TSP experiments were repeated ten times. We obtained the ground truth for each snapshot by formulating it as a QUBO problem and solving it using the Gurobi solver (Gurobi Optimization, LLC, 2024) with a time limit of 60 seconds. Additional implementation details are provided in Appendix C.
@@ 5.3 Empirical results on DCO problems
Tables 1, 2, and 3 report the mean ApR of all methods on dynamic MaxCut, MIS, and TSP, respectively. We also evaluate variants of DyCO-GNN that apply SP to different layers of DyCO-GNN. For MaxCut and MIS, results are reported using the final model checkpoint. For TSP, both the final and the best-performing checkpoints are considered. The wall-clock time of the best-performing checkpoint includes the total decoding time over all checkpoints available. As noted in Section 4, the advantage of warm-started PI-GNN compared to static PI-GNN diminishes quickly when we relax the runtime constraint. DyCO-GNN closes this performance gap and consistently outperforms static PI-GNN and warm-started PI-GNN across all problems and datasets. Even on a strict budget (first column of each table), DyCO-GNN surpasses warm-started PI-GNN in most cases. Crucially, DyCO-GNN consistently finds better solutions than fully converged static PI-GNN, often within just 1.67% to 33.33% of the total runtime. This improvement possibly stems from information carryover and continual learning on a similar graph structure, enabled by the utilization of the previously learned hypothesis. Although the three versions of DyCO-GNN give strong results, no single setting dominates all tasks and datasets. We will discuss the implications in Section 6. We visualize snapshot-level ApRs under different time budgets in Figures 2, 3, and 4. Notably, DyCO-GNN maintains its edge across almost all evaluated snapshots, confirming its robustness. Additional experimental results are provided in Appendix D.
@@
Method 500/0.75 1000/1.44 2000/2.81 3000/4.18 5000/6.98 10000/13.95 20000/27.84
st70 PI-GNN (static) 1.19948 1.19016 1.15768 1.14324 1.11844 1.09454 1.07768
PI-GNN (warm) 1.18218 1.16969 1.18631 1.17347 1.14945 1.14242 1.11914
DyCO-GNN (emb) 1.15915 1.12257 1.10795 1.10227 1.09809 1.07375 1.07743
DyCO-GNN (GNN) 1.16442 1.12976 1.08656 1.08362 1.07243 1.07030 1.05035
DyCO-GNN (full) 1.14632 1.11718 1.07749 1.06831 1.06763 1.04521 1.04617
Sensitivity of different methods to varying degrees of change. Tables 1 and 2 show that the quality gap of converged solutions between static PI-GNN and DyCO-GNN (resp. warm-started PI-GNN) is larger (resp. smaller) for DBLP. This is because we constructed the DTDG for DBLP with a smaller relative change in the edge set ($\sim$2% of total edges), which leads to greater structural overlaps. Consequently, warm-started PI-GNN and DyCO-GNN would be stronger in such a scenario. We further conducted a sensitivity analysis by varying the degrees of change in the DTDGs. Results are illustrated in Figure 5. When $\Delta\text{edges}$ is getting smaller, the gap between static and warm-started PI-GNN is narrowed, and the advantage of DyCO-GNN compared to static PI-GNN is more salient, which is consistent with our hypothesis.
@@
Sensitivity analysis on SP parameters. To demonstrate the generality and robustness of our method without relying on dataset-specific tuning, we deliberately chose a single, reasonable set of SP parameter values (i.e., $\lambda_{\text{shrink}}$ and $\lambda_{\text{perturb}}$) and applied it uniformly, highlighting that strong performance can be achieved without sensitive hyperparameter tuning. For completeness, we include a sensitivity analysis on the SP parameters in Appendix D.4.
@@ 6 Discussion
Conclusion. We introduced DyCO-GNN, the first general UL-based framework designed to solve DCO problems. Through extensive experiments on dynamic MaxCut, MIS, and TSP, we demonstrated that DyCO-GNN consistently outperforms both static and warm-started PI-GNN baselines under varying runtime constraints and across diverse problem settings. Our analysis reveals that DyCO-GNN achieves superior solution quality even under strict time budgets, always surpassing the best-performing checkpoints of the baseline methods in a fraction of their runtime. These improvements are significant in rapidly changing environments, where timely decision-making is critical. Furthermore, DyCO-GNN’s ability to efficiently adapt to evolving problem instance snapshots without optimization from scratch or external supervision underscores its potential for real-world deployment in dynamic, resource-constrained settings.
Limitations and future work. We have shown that applying SP to different layers of DyCO-GNN yields varying performance. Currently, the embedding layer and the GNN layers are jointly optimized without explicit decoupling. By introducing appropriate regularization strategies or auxiliary loss functions, we can promote distinct functional roles across layers; for instance, encouraging the embedding layer to capture representations specific to both the problem and graph structure and the GNN layers to learn to aggregate such representations. Additionally, a promising direction for future work is to make the SP step adaptive. If we evaluate the likelihood of changes in node assignments under dynamic conditions based on certain heuristics, it becomes possible to apply SP in a node-wise, adaptive manner.
@@ Appendix A Additional related work
DCO problems have been addressed from a theoretical perspective (Onak and Rubinfeld, 2010; Thorup, 2007), including for the dynamic MaxCut (Wasim and King, 2020), MIS (Assadi et al., 2018), and TSP (Ausiello et al., 2009) problems, as well as from the perspective of designing heuristics (Yang et al., 2012). Our work brings a learning-based approach to this line of work.
The idea of warm-starting semidefinite programs in general, and the starting solutions of interior-point methods (which can be used to solved SDPs, among other problem classes) has been explored from a theoretical perspective (Yildirim and Wright, 2002; John and Yıldırım, 2008; Angell and Mccallum, 2024). Our theoretical findings complement this literature by exploring the impacts of warm-starting SDP solutions for the MaxCut problem. Our learning-based method DyCO-GNN also brings similar ideas into the realm of learning-based DCO methods.
Appendix B Dataset statistics
Original Preprocessed
Dataset Nodes Edges Nodes Edges $\Delta$ edges
Infectious 410 17,298 410 2765 $\sim$276 (10%)
UC Social 1899 59,835 1899 13,838 $\sim$1384 (10%)
DBLP 12,590 49,759 12,590 49,636 $\sim$993 (2%)
burma14 14 91 15 105 -
ulysses22 22 231 23 253 -
st70 70 2,415 71 2,485 -
Appendix C Additional implementation details
We used the Adam optimizer (Kingma and Ba, 2015); the learning rate was set to 0.001 for all MaxCut and MIS experiments, 0.0002 for burma14 and st70, and 0.0005 for ulysses22. As described in Algorithms 1 and 2, we optimize over the first snapshot for $\text{epoch}_{\text{max}}$ epochs. Thus, we skip the first snapshot during evaluation since the results will be identical across all methods. We set $\text{epoch}_{\text{max}}=3000$ for all MaxCut and MIS experiments, $\text{epoch}_{\text{max}}=10000$ for burma14 and ulysses22, and $\text{epoch}_{\text{max}}=20000$ for st70.
The penalty term $M$ was set to 2 for MIS and $2\times\max_{(i,j)\in E}w_{ij}$ (i.e., $2\timesthe$ max distance between any two nodes) for TSP. For MIS, we postprocessed the output of our model by greedily removing violations. For TSP, we decoded the route step by step. More specifically, for burma14 and ulysses22, we discarded nodes that have already been visited and selected the most likely node to visit based on our model output; for st70, we found that this greedy decoding failed to find good enough routes and instead applied beam search that expands the top 20 possible valid nodes at each step.
All models were implemented using PyTorch (Paszke et al., 2019) and PyTorch Geometric (Fey and Lenssen, 2019). Experiments were conducted on a machine with a single NVIDIA GeForce RTX 4090 GPU, a 32-core Intel Core i9-14900K CPU, and 64 GB of RAM running Ubuntu 24.04.
Appendix D Additional experimental results
D.1 DyCO-GNN with SAGEConv for dynamic MaxCut and MIS